OpenStudy (anonymous):

Let f(x) = (x + 6) / (x - 10) and g(x) = sqrt( x+1) . Find the domains of the following composite functions. Enter your answer in interval notation. Domain of (f\circ g)(x) = is

OpenStudy (anonymous):

Domain of (f\circ g)(x) = is also means: (f o g)(x) ...

OpenStudy (anonymous):

(f o g)(x) = f(g(x)) = f(sqrt(x+1)) = (sqrt(x+1) + 6)/(sqrt(x+1) - 10)

OpenStudy (anonymous):

\[\large (f \circ g)(x)=\frac{\sqrt{x+1}+6}{\sqrt{x+1}-10}\]

OpenStudy (anonymous):

dom(f o g) = [-1,inf) , exclude x=99, 10